A family has five children. The probability of having a girl is 1/2. What is the probability of having no more than 3 boys?
Answer #1

Let p=P(boy) = 1-P(girl) = 1-(1/2) = 1/2 = 0.5

Let X denote the number of boys.

Then X follows a Binomial distribution with n=5 and p=0.5.

P(x) = nCxp*(1-p)*-* = 5C;(0.5)*(0.5)5-* = 5C(0.5). (r = 0,1,2,3,4,5) P(no more than 3 boys) = P(X 53) = P(0)+ P(1)+P(2)+P(3